### v3

# IRIS: LLM-Assisted Static Analysis for Detecting Security Vulnerabilities

Abstract
Software is prone to security vulnerabilities. Program analysis tools to detect them have limited effectiveness in practice due to their reliance on human labeled specifications. Large language models (or LLMs) have shown impressive code generation capabilities but they cannot do complex reasoning over code to detect such vulnerabilities especially since this task requires whole-repository analysis. We propose IRIS, a neuro-symbolic approach that systematically combines LLMs with static analysis to perform whole-repository reasoning for security vulnerability detection. Specifically, IRIS leverages LLMs to infer taint specifications and perform contextual analysis, alleviating needs for human specifications and inspection. For evaluation, we curate a new dataset, CWE-Bench-Java, comprising 120 manually validated security vulnerabilities in real-world Java projects. A state-of-the-art static analysis tool CodeQL detects only 27 of these vulnerabilities whereas IRIS with GPT-4 detects 55 ($+28$) and improves upon CodeQL’s average false discovery rate by 5% points. Furthermore, IRIS identifies 4 previously unknown vulnerabilities which cannot be found by existing tools. IRIS is available publicly at https://github.com/iris-sast/iris.
1 Introduction
Security vulnerabilities pose a major threat to the safety of software applications and its users. In 2023 alone, more than 29,000 CVEs were reported—almost 4000 higher than in 2022 (CVE Trends, ). Detecting vulnerabilities is extremely challenging despite advances in techniques to uncover them. A promising such technique called static taint analysis is widely used in popular tools such as GitHub CodeQL (Avgustinov et al., 2016), Facebook Infer (FB Infer, ), Checker Framework (Checker Framework, ), and Snyk Code (Snyk.io, ). These tools, however, face several challenges that greatly limit their effectiveness and accessibility in practice.
Figure 1: Overview of the IRIS neuro-symbolic system. It checks a given whole repository for a given type of vulnerability (CWE) and outputs a set of potential vulnerable paths with explanations.
False negatives due to missing taint specifications of third-party library APIs. First, static taint analysis predominantly relies on specifications of third-party library APIs as sources, sinks, or sanitizers. In practice, developers and analysis engineers have to manually craft such specifications based on their domain knowledge and API documentation. This is a laborious and error-prone process that often leads to missing specifications and incomplete analysis of vulnerabilities. Further, even if such specifications may exist for many libraries, they need to be periodically updated to capture changes in newer versions of such libraries and also cover new libraries that are developed.
False positives due to lack of precise context-sensitive and intuitive reasoning. Second, it is well-known that static analysis often suffers from low precision, i.e., it may generate many false alarms (Kang et al., 2022; Johnson et al., 2013). Such imprecision stems from multiple sources. For instance, the source or sink specifications may be spurious, or the analysis may over-approximate over branches in code or possible inputs. Further, even if the specifications are correct, the context in which the detected source or sink is used may not be exploitable. Hence, a developer may need to triage through several potentially false security alerts, wasting significant time and effort.
Limitations of prior data-driven approaches to improve static taint analysis. Many techniques have been proposed to address the challenges of static taint analysis. For instance, Livshits et al. (2009) proposed a probabilistic approach, MERLIN, to automatically mine taint specifications. A more recent work, Seldon (Chibotaru et al., 2019), improves the scalability of this approach by formulating the taint specification inference problem as a linear optimization task. However, such approaches rely on analyzing the code of third-party libraries to extract specifications, which is expensive and hard to scale. Researchers have also developed statistical and learning-based techniques to mitigate false positive alerts (Jung et al., 2005; Heckman & Williams, 2009; Hanam et al., 2014). However, such approaches still have limited effectiveness in practice (Kang et al., 2022).
Large Language Models (or LLMs) have made impressive strides in code generation and summarization. LLMs have also been applied to code related tasks such as program repair (Xia et al., 2023), code translation (Pan et al., 2024), test generation (Lemieux et al., 2023), and static analysis (Li et al., 2024). Recent studies (Steenhoek et al., 2024; Khare et al., 2023) evaluated LLMs’ effectiveness at detecting vulnerabilities at the method level and showed that LLMs fail to do complex reasoning with code, especially because it depends on the context in which the method is used in the project. On the other hand, recent benchmarks like SWE-Bench (Jimenez et al., 2023) show that LLMs are also poor at doing project-level reasoning. Hence, an intriguing question is whether LLMs can be combined with static analysis to improve their reasoning capabilities. In this work, we answer this question in the context of vulnerability detection and make the following contributions:
Approach. We propose IRIS, a neuro-symbolic approach for vulnerability detection that combines the strengths of static analysis and LLMs. Fig. 1 presents an overview of IRIS. Given a project to analyze for a given vulnerability class (or CWE), IRIS applies LLMs for mining CWE-specific taint specifications. IRIS augments such specifications with CodeQL, a tool for static taint analysis. Our intuition here is because LLMs have seen numerous usages of such library APIs, they have an understanding of the relevant APIs for different CWEs. Further, to address the imprecision problem of static analysis, we propose a contextual analysis technique with LLMs that reduces the false positive alarms and minimizes the triaging effort for developers. Our key insight is that encoding the code-context and path-sensitive information in the prompt elicits more reliable reasoning from LLMs. Finally, our neuro-symbolic approach allows LLMs to do more precise whole-repository reasoning and minimizes the human effort involved in using static analysis tools.
Dataset. We curate a dataset of manually vetted and compilable Java projects, CWE-Bench-Java, containing 120 vulnerabilities (one per project) across four common vulnerability classes. The projects in the dataset are complex, containing 300K lines of code on average, and 10 projects with more than a million lines of code each, making it a challenging benchmark for vulnerability detection. The dataset and the corresponding scripts to fetch, build, and analyze the Java projects are available publicly at https://github.com/iris-sast/cwe-bench-java.
Results. We evaluate IRIS on CWE-Bench-Java using 7 diverse open- and closed-source LLMs. Overall, IRIS obtains the best results with GPT-4, detecting 55 vulnerabilities, which is 28 ($103.7\%$) more than CodeQL, the existing best-performing static analyzer. We show that the increase is not at the expense of false positives, as IRIS with GPT-4 achieves an average false discovery rate of $84.82\%$, which is $5.21\%$ lower than that of CodeQL. Further, when applied to the latest versions of 30 Java projects, IRIS with GPT-4 discovered 4 previously unknown vulnerabilities.
2 Motivating Example
Figure 2: An example of Code Injection (CWE-94) vulnerability found in cron-utils (CVE-2021-41269) that CodeQL fails to detect. We number the program points of the vulnerable path.
We illustrate the effectiveness of IRIS in detecting a previously known code-injection (CWE-094) vulnerability in cron-utils (ver. 9.1.5), a Java library for Cron data manipulation. Fig. 2 shows the relevant code snippets. A user-controlled string value passed into isValid function is transferred without sanitization to the parse function. If an exception is thrown, the function constructs an error message with the input. However, the error message is used to invoke method buildConstraintViolationWithTemplate of class ConstraintValidatorContext in javax.validator, which interprets the message string as a Java Expression Language (Java EL) expression. A malicious user may exploit this vulnerability by crafting a string containing a shell command such as Runtime.exec(‘rm -rf /’) to delete critical files on the server.
Detecting this vulnerability poses several challenges. First, the cron-utils library consists of 13K SLOC (lines of code excluding blanks and comments), which needs to be analyzed to find this vulnerability. This process requires analyzing data and control flow across several internal methods and third-party APIs. Second, the analysis needs to identify relevant sources and sinks. In this case, the value parameter of the public isValid method may contain arbitrary strings when invoked, and hence may be a source of malicious data. Additionally, external APIs like buildConstraintViolationWithTemplate can execute arbitrary Java EL expressions, hence they should be treated as sinks that are vulnerable to Code Injection attacks. Finally, the analysis also requires identifying any sanitizers that block the flow of untrusted data.
Modern static analysis tools, like CodeQL, are effective at tracing taint data flows across complex codebases. However, CodeQL fails to detect this vulnerability due to missing specifications. CodeQL includes many manually curated specifications for sources and sinks across more than 360 popular Java library modules. However, manually obtaining such specifications requires significant human effort to analyze, specify, and validate. Further, even with perfect specifications, CodeQL may often generate numerous false positives due to a lack of contextual reasoning, increasing the developer’s burden of triaging the results.
In contrast, IRIS takes a different approach by inferring project- and vulnerability-specific specifications on-the-fly by using LLMs. The LLM-based components in IRIS correctly identify the untrusted source and the vulnerable sink. IRIS augments CodeQL with these specifications and successfully detects the unsanitized dataflow path between the detected source and sink in the repository. However, augmented CodeQL produces many false positives, which are hard to eliminate using logical rules. To solve this challenge, IRIS encodes the detected code paths and the surrounding context into a simple prompt and uses an LLM to classify it as true or false positive. Specifically, out of 8 paths reported by static analysis, 5 false positives are filtered out, leaving the path in Fig. 2 as one of the final alarms. Overall, we observe that IRIS can detect many such vulnerabilities that are beyond the reach of CodeQL-like static analysis tools, while keeping false alarms to a minimum.
3 IRIS Framework
At a high level, IRIS takes a Java project $P$, the vulnerability class $C$ to detect, and a large language model LLM, as inputs. IRIS statically analyzes the project $P$, checks for vulnerabilities specific to $C$, and returns a set of potential security alerts $A$. Each alert is accompanied by a unique code path from a taint source to a taint sink that is vulnerable to $C$ (i.e., the path is unsanitized).
As illustrated in Fig. 3, IRIS has four main stages: First, IRIS builds the given Java project and uses static analysis to extract all candidate APIs, including invoked external APIs and internal function parameters. Second, IRIS queries an LLM to label these APIs as sources or sinks that are specific to the given vulnerability class $C$. Third, IRIS transforms the labeled sources and sinks into specifications that can be fed into a static analysis engine, such as CodeQL, and runs a vulnerability class-specific taint analysis query to detect vulnerabilities of that class in the project. This step generates a set of vulnerable code paths (or alerts) in the project. Finally, IRIS triages the generated alerts by automatically filtering false positives, and presents them to the developer.
Figure 3: An illustration of the IRIS pipeline.
3.1 Problem Statement
We formally define the static taint analysis problem for vulnerability detection. Given a project $P$, taint analysis extracts an inter-procedural data flow graph $\mathbb{G}=(\mathbb{V},\mathbb{E})$, where $\mathbb{V}$ is the set of nodes representing program expressions and statements, and $\mathbb{E}\subseteq\mathbb{V}\times\mathbb{V}$ is the set of edges representing data or control flow edges between the nodes. A vulnerability detection task comes with two sets $\bm{V}^{C}_{\textit{source}}\subseteq\mathbb{V}$, $\bm{V}^{C}_{\textit{sink}}\subseteq\mathbb{V}$ that denote source nodes where tainted data may originate and sink nodes where a security vulnerability can occur if tainted data reaches it, respectively. Naturally, different classes $C$ of vulnerabilities (or CWEs) have different source and sink specifications. Additionally, there can be sanitizer specifications, $\bm{V}^{C}_{\textit{sanitizer}}\in\mathbb{V}$, that block the flow of tainted data (such as escaping special characters in strings).
[…]
3.2 Candidate Source/Sink Extraction
[…]
To identify taint specifications $\bm{S}_{\textit{source}}^{C}$ and $\bm{S}_{\textit{sink}}^{C}$, we first extract $\bm{S}^{\text{ext}}$: external library APIs that are invoked in the given Java project and are potential candidates to be taint sources or sinks. We also extract $\bm{S}^{\text{int}}$, internal library APIs that are public and may be invoked by a downstream library. We use CodeQL to extract such candidates and their corresponding metadata such as method name, type signature, enclosing packages and classes, and even JavaDoc documentations, if applicable.
3.3 Inferring Taint Specifications using LLMs
We develop an automated specification inference technique: $\textit{LabelSpecs}(\bm{S}^{\#},\texttt{LLM},C,R)=\bm{S}_{R}^{C}$, where $\bm{S}^{\#}=\bm{S}^{\text{ext}}\cup\bm{S}^{\text{int}}$ are candidate specifications for sources and sinks. In this work, we do not consider sanitizer specifications, because they typically do not vary for the vulnerability classes that we consider. We use LLMs to infer taint specifications. Specifically, external APIs in $\bm{S}^{\text{ext}}$ can be classified as either source or sink, while internal APIs in $\bm{S}^{\text{int}}$ can have their formal parameters identified as sources. In the Appendix, we show the user prompts for inferring source and sink specifications from external APIs and internal function formal parameters.
Due to the sheer number of APIs to be labeled, we insert a batch of APIs in a single prompt and ask the LLM to respond with JSON formatted strings. The batch size is a tunable hyper-parameter. We adopt few-shot (usually 3-shot) prompting strategy for labeling external APIs, while zero-shot is used for labeling internal APIs. Notably for internal APIs, we also include information from repository readme and JavaDoc documentations, if applicable. In practice, we find that this extra information helps LLM understand the high-level purpose and usage of the codebase, resulting in better labeling accuracy. At the end of this stage, we have successfully obtained $\bm{S}_{\textit{source}}^{C}$ and $\bm{S}_{\textit{sink}}^{C}$ which are going to be used by the static analysis engine in the next stage.
Figure 4: LLM user prompt and response for contextual analysis of dataflow paths. In the user prompt, we mark with color the CWE and path information that is filling the prompt template. For cleaner presentation, we modify the snippets and left out the system prompt.
3.4 Vulnerability Detection
[…]
$\textit{CodeQL}(\mathbb{G}^{P},\bm{S}_{\textit{source}}^{C},\bm{S}_{\textit{sink}}^{C},\textit{Query}^{C})=\{\textit{Path}_{1},\ldots,\textit{Path}_{k}\}.$
CodeQL itself contains numerous specifications of third-party APIs for each vulnerability class. However, as we show later in our evaluation, despite having such specialized queries and extensive specifications, CodeQL fails to detect a majority of vulnerabilities in real-world projects. For our analysis, we write a specialized CodeQL query for each vulnerability that uses our mined specifications instead of those provided by CodeQL. Details of our queries are described in the Appendix.
3.5 Triaging of Alerts via Contextual Analysis
[…]
Fig. 4 presents an example prompt for contextual analysis. The prompt includes CWE information and code snippets for nodes along the path, with an emphasis on the source and sink. For the intermediate steps, we include the file names and the line of code. When the path is too long, we keep only a subset of nodes to limit the size of the prompt. More details and design decisions of this prompt are described in the Appendix. We expect the LLM to respond in JSON format with the final verdict as well as an explanation to the verdict. The JSON format prompts the LLM to generate the explanation before delivering the final verdict, as presenting the judgment after the reasoning process is known to yield better results. In addition, if the verdict is false, we ask the LLM to indicate whether the source or sink is a false positive, which helps to prune other paths and thereby save on the number of calls to the LLM.
3.6 Evaluation Metrics
We evaluate the performance of IRIS and its baselines using three key metrics: number of vulnerability detected (#Detected), average false discovery rate (AvgFDR), and average F1 (AvgF1). For evaluation, we assume that we have a dataset $\mathcal{D}=\{P_{1},\dots,P_{n}\}$ where each $P_{i}$ is a Java project, and known to contain at least one vulnerability. The label for a project $P$ is provided as a set of crucial program points $\mathbf{V}_{\text{vul}}^{P}=\{V_{1},\dots,V_{n}\}$ where the vulnerable paths should pass through, indicated by $\textit{Path}\cap\mathbf{V}_{\text{vul}}^{P}\neq\emptyset$. In practice, these are typically the patched methods that can be collected from each vulnerability report. If at least one detected vulnerable path passes through a fixed location for the given vulnerability, then we consider the vulnerability detected. Let $\textit{Paths}^{P}$ be the set of detected paths for each project $P$ from prior stages. The metrics are formally defined as follows:
#$\begin{array}[]{rlrl}\textit{\#VulPath}(P)=&|\{\textit{Path}\in\textit{Paths}^{P}~{}|~{}\textit{Path}~{}\cap~{}\mathbf{V}_{\text{vul}}^{P}\neq\emptyset\}|,&\textit{Rec}(P)=&\mathbbm{1}_{\textit{\#VulPath}(P)>0},\\ \textit{\#Detected}(\mathcal{D})=&\textstyle\sum_{P\in\mathcal{D}}\textit{Rec}(P),&\textit{Prec}(P)=&\textstyle\frac{\textit{\#VulPath}(P)}{|\textit{Paths}^{P}|},\\ \textit{AvgFDR}(\mathcal{D})=&\text{avg}_{P\in\mathcal{D},|\textit{Paths}^{P}|>0}1-\textit{Prec}(P),&\textit{AvgF1}(\mathcal{D})=&\textstyle\frac{1}{|\mathcal{D}|}\textstyle\sum_{P\in\mathcal{D}}\frac{2\cdot\textit{Prec}(P)\cdot\textit{Rec}(P)}{\textit{Prec}(P)+\textit{Rec}(P)}.\end{array}$
Specifically, a lower AvgFDR is preferable, as it indicates a lower ratio of false positives. We note that $\textit{Prec}(P)$ might sometimes be undefined due to division-by-zero if the detection tool retrieves no path ($|\textit{Paths}^{P}|=0$). Therefore, for AvgFDR to be meaningful, we only consider the projects where at least one positive result is produced ($|\textit{Paths}^{P}|>0$). AvgF1 avoids this issue since $\textit{Rec}(P)=0$ when no positive labels exist, forcing the F1 term to be zero regardless of $\textit{Prec}(P)$.
4 CWE-Bench-Java: A Dataset of Security Vulnerabilities in Java
To evaluate our approach, we require a dataset of vulnerable versions of Java projects with several important characteristics: 1) Each benchmark should have relevant vulnerability metadata, such as the CWE ID, CVE ID, fix commit, and vulnerable project version, 2) each project in the dataset must be compilable, which is a key requirement for static analysis and data flow graph extraction, 3) the projects must be real-world, which are typically more complex and hence challenging to analyze compared to synthetic benchmarks, and 4) finally, each vulnerability and its location (e.g., method) in the project must be validated so that this information can be used for robust evaluation of vulnerability detection tools. Unfortunately, no existing dataset satisfies all these requirements.
Figure 5: Steps for curating CWE-Bench-Java, and dataset statistics.
To address these requirements, we curate our own dataset of vulnerabilities. For this paper, we focus only on vulnerabilities in Java libraries that are available via the widely used Maven package manager. We choose Java because it is commonly used to develop server-side, Android, and web applications, which are prone to security risks. Further, due to Java’s long history, there are many existing CVEs in numerous Java projects that are available for analysis. We initially use the GitHub Advisory database (GitHub, 2024b; c) to obtain such vulnerabilities, and further filter it with cross-validated information from multiple sources, including manual verification. Fig. 5 illustrates the complete set of steps for curating CWE-Bench-Java.
As shown in the statistics (Fig. 5), the sheer size of these projects make them challenging to analyze for any static analysis tool or ML-based tool. Each project in CWE-Bench-Java comes with GitHub information, vulnerable and fix version, CVE metadata, a script that automatically fetches and builds, and the set of program locations that involve the vulnerability.
5 Evaluation
We perform extensive experimental evaluations of IRIS and demonstrate its practical effectiveness in detecting vulnerabilities in real-world Java repositories in CWE-Bench-Java. While additional results and analyses are provided in the Appendix, we address the following key research questions:
  • RQ 1: How many previously known vulnerabilities can IRIS detect?
  • RQ 2: Does IRIS detect new, previously unknown vulnerabilities?
  • RQ 3: How good are the inferred source/sink specifications by IRIS?
  • RQ 4: How effective are the individual components of IRIS?
5.1 Experimental Setup
We select two closed-source LLMs from OpenAI: GPT-4 (gpt-4-0125-preview) and GPT-3.5 (gpt-3.5-turbo-0125) for our evaluation. We also select instruction-tuned versions of four open-source LLMs via huggingface API: Llama 3 (L3) 8B and 70B, Qwen-2.5-Coder (Q2.5C) 32B, Gemma-2 (G2) 27B, and DeepSeekCoder (DSC) 7B. For the CodeQL baseline, we use version 2.15.3 and its built-in Security queries specifically designed for each CWE. Other baselines included are Facebook Infer (FB Infer, ), SpotBugs (Lavazza et al., 2020), and Snyk (Snyk.io, ). We expand further on the other experimental setups in the Appendix.
5.2 RQ1: Effectiveness of IRIS on Detecting Existing Vulnerabilities
Table 1: Overall performance comparison of CodeQL vs IRIS on Detection Rate ($\uparrow$), Average FDR ($\downarrow$), and Average F1 ($\uparrow$). We present results of IRIS with LLMs including GPT-4 and GPT-3.5, L3 8B and 70B, Q2.5C 32B, G2 27B, and DSC 7B.
[…]
CodeQL $27$ $22.50$ $90.03$ $0.076$
IRIS + GPT-4 $\mathbf{55}$ ($\uparrow 28$) $\mathbf{45.83}$ ($\uparrow 23.33$) $\mathbf{84.82}$ ($\downarrow 5.21$) $\mathbf{0.177}$ ($\uparrow 0.101$)
GPT-3.5 $47$ ($\uparrow 20$) $39.17$ ($\uparrow 16.67$) $90.42$ ($\uparrow 0.39$) $0.096$ ($\uparrow 0.020$)
L3 8B $41$ ($\uparrow 14$) $34.17$ ($\uparrow 11.67$) $95.55$ ($\uparrow 5.52$) $0.058$ ($\downarrow 0.018$)
L3 70B $54$ ($\uparrow 27$) $45.00$ ($\uparrow 22.50$) $90.96$ ($\uparrow 0.93$) $0.113$ ($\uparrow 0.037$)
Q2.5C 32B $47$ ($\uparrow 20$) $39.17$ ($\uparrow 16.67$) $92.38$ ($\uparrow 2.35$) $0.097$ ($\uparrow 0.021$)
G2 27B $45$ ($\uparrow 18$) $37.50$ ($\uparrow 15.00$) $91.23$ ($\uparrow 1.20$) $0.100$ ($\uparrow 0.024$)
DSC 7B $52$ ($\uparrow 25$) $43.33$ ($\uparrow 20.83$) $95.40$ ($\uparrow 5.37$) $0.062$ ($\downarrow 0.014$)
The results in Table 1 highlight IRIS’s superior performance compared to CodeQL. Specifically, IRIS, when paired with GPT-4, identifies 55 vulnerabilities—28 more than CodeQL. While GPT-4 shows the highest efficacy, smaller, specialized LLMs like DeepSeekCoder 7B still detect 52 vulnerabilities, suggesting that our approach can effectively leverage smaller-scale models, enhancing accessibility. Notably, this increase in detected vulnerabilities does not compromise precision, as evidenced by IRIS’s lower average false discovery rate (FDR) with GPT-4 compared to CodeQL. Moreover, IRIS improves average F1 by 0.1, reflecting a better balance between precision and recall. We note that the reported average FDR is a coarse measure as our metrics may consider a true (but unknown) vulnerability found by IRIS as a false positive. Hence, the reported FDR is an upper bound. To get a better sense of detection accuracy, we manually analyzed 50 random alarms reported by IRIS (using GPT-4) and found that 27/50 alarms exhibit potential attack surfaces, yielding a more refined estimated false discovery rate of 46%. Hence, IRIS will likely be more effective in practice.
Table 2 presents a detailed breakdown of detected vulnerabilities, comparing IRIS against various baselines. With the exception of IRIS using Llama-3 8B, which underperforms in detecting CWE-22 vulnerabilities, IRIS consistently outperforms all other baselines. Notably, CWE-78 (OS Command Injection) remains particularly challenging for all LLMs. Our manual investigation revealed that the vulnerability patterns in CWE-78 are highly intricate, often involving OS command injections via gadget-chains (Cao et al., 2023) or external side effects, such as file writes, which are difficult to track using static analysis. This highlights the inherent limitations of static analysis, as opposed to dynamic approaches—an area that we leave for future work.
Table 2: Per-CWE statistics of number of vulnerabilities detected (#Detected) by baselines and IRIS. The compared baselines are CodeQL (QL), Facebook Infer (Infer), Spotbugs (SB), and Snyk. The values in parentheses show the differences of detection by IRIS against CodeQL.
CWE #Vuls Baselines IRIS with
QL Infer SB Snyk GPT-4 GPT-3.5 L3 8B L3 70B DSC 7B
CWE-22 55 22 0 2 21 31 ($\uparrow 9$) 25 ($\uparrow 3$) 19 ($\downarrow 3$) 29 ($\uparrow 7$) 25 ($\uparrow 3$)
CWE-78 13 1 0 1 1 3 ($\uparrow 2$) 1 ($=0$) 2 ($\uparrow 1$) 2 ($\uparrow 1$) 3 ($\uparrow 2$)
CWE-79 31 4 0 1 1 13 ($\uparrow 9$) 13 ($\uparrow 9$) 9 ($\uparrow 9$) 14 ($\uparrow 10$) 14 ($\uparrow 10$)
CWE-94 21 0 0 0 0 8 ($\uparrow 8$) 8 ($\uparrow 8$) 11 ($\uparrow 11$) 9 ($\uparrow 9$) 10 ($\uparrow 10$)
All 120 27 0 4 23 55 ($\uparrow 28$) 47 ($\uparrow 20$) 41 ($\uparrow 14$) 54 ($\uparrow 27$) 52 ($\uparrow 25$)
5.3 RQ2: Previously Unknown Vulnerabilities by IRIS
We applied IRIS with GPT-4 to the latest versions of 30 Java projects. Among the 16 inspected projects where IRIS raised at least one alert, we identified 4 vulnerabilities, including 3 instances of path injection (CWE-22) and one case of code-injection (CWE-94). To ensure that these vulnerabilities were indeed uncovered due to IRIS’s integration with LLMs, we verified that CodeQL alone did not detect them. We highlight one such vulnerability in Fig. 8. CodeQL was unable to detect this issue due to a missing source specification, while GPT-4 successfully flagged the API endpoint restoreFromCheckpoint as a potential entry point for attack.
5.4 RQ3: Quality of LLM-Inferred Taint Specifications
Figure 6: Recall of LLM-inferred taint specifications against CodeQL’s taint specifications.
Figure 7: Estimated precision of LLM-inferred specifications on randomly sampled labels.
Figure 8: A previously unknown vulnerability found in alluxio 2.9.4. The snippets are slightly modified for presentation purpose. A user with database restoration permission may supply a database checkpoint Zip file with malicious entry name. When unzipped, the entry may be written to an arbitrary directory, causing a Zip-Slip vulnerability (CWE-022) that could corrupt the hosting server.
The LLM-inferred taint specifications are fundamental to IRIS’s effectiveness. To assess the quality of these specifications, we conducted two experiments. First, we used CodeQL’s taint specifications as a benchmark to estimate the recall of both source and sink specifications inferred by LLMs (Fig. 6). However, since CodeQL offers a limited set of specifications, we also needed to assess the quality of inferred specifications outside of its known coverage. To this end, we manually analyzed 960 randomly selected samples of LLM-inferred source and sink labels (30 per combination of CWE and LLM) and estimated the overall precision of the specifications (Fig. 7).
LLM-inferred sinks can replace CodeQL sinks. Overall, LLMs demonstrated high recall when tested against CodeQL’s sink specifications (Fig. 6), with GPT-4 scoring the highest ($87.11\%$). While the recall for source specifications was generally lower, we found that CodeQL tends to over-approximate its source specifications to compensate for a low detection rate. On the other hand, GPT-4 achieved high precision (over 70%) in manual evaluations (Fig. 7), aligning with the lower false discovery rate previously reported in Table 1. For other LLMs, the combination of high recall but lower precision suggests a tendency to over-approximate sink specifications.
Over-approximating specifications can benefit IRIS. Although the precision for LLMs other than GPT-4 is lower, over-approximation can actually help address a core limitation of CodeQL—its restricted set of taint specifications. By over-approximating, LLMs expand the coverage of taint analysis, offering a partial solution to CodeQL’s limited scope. The impact of this imprecision can be mitigated through contextual analysis as we show next in the ablation studies.
5.5 RQ4: Ablation Studies
Both LLM-inferred sources and sinks are necessary. Table 3 presents additional results when using either only the source or sink specification from an LLM in IRIS. For this experiment, we only use the results with GPT-4 for comparison. Each row present the number of detected vulnerabilities per CWE. We observe that omitting either source or sink specifications inferred by GPT-4 causes a drastic reduction in overall recall.
Performance gain of contextual analysis depends on LLM’s reasoning capability. As shown in Fig. 9, contextual analysis is highly necessary for the precision and F1 score improvements. However, only GPT-4, GPT-3.5, and Llama-3 70B see a positive impact after contextual analysis, while the smaller models see negative. The false positive reduction of contextual analysis is the most effective when the LLM possesses decent reasoning capability. Indeed, smaller models are more likely to respond with “vulnerable” than larger models.
Table 3: Ablation on LLM inferred source and sink specifications (CodeQL (QL) versus GPT-4), evaluated using the #Detected metrics. When replacing either source or sink with CodeQL specs, we see significantly less vulnerabilities detected.
CWE 22 78 79 94 Total
$\text{Src}_{\text{QL}}$ $+$ $\text{Snk}_{\text{QL}}$ 22 1 4 0 27 ($\downarrow 28$)
$\text{Src}_{\text{GPT4}}$ $+$ $\text{Snk}_{\text{QL}}$ 28 3 5 0 36 ($\downarrow 19$)
$\text{Src}_{\text{QL}}$ $+$ $\text{Snk}_{\text{GPT4}}$ 10 1 9 4 24 ($\downarrow 31$)
$\text{Src}_{\text{GPT4}}$ $+$ $\text{Snk}_{\text{GPT4}}$ 31 3 13 8 55
Figure 9: Improvements of Avg. Precision and Avg. F1 after contextual analysis.
6 Related Work
Learning-based approaches for vulnerability detection. Numerous prior techniques incorporate deep learning for detecting vulnerabilities. This includes Graph Neural Network based models such as Zhou et al. (2019); Chakraborty et al. (2020); Dinella et al. (2020); Hin et al. (2022); Li et al. (2021a); LSTM-based models such as Li et al. (2020; 2021b); and fine-tuning of Transformer-based models such as Fu & Tantithamthavorn (2022); Steenhoek et al. (2023); Cheng et al. (2022). These approaches focus on method-level detection of vulnerabilities and provide only a binary label classifying a method as vulnerable or not. In contrast, IRIS performs whole-project analysis and provides a distinct code path from a source to a sink and can be tailored for detecting different CWEs. More recently, multiple studies demonstrated that LLMs are not effective at detecting vulnerabilities in real-world code (Steenhoek et al., 2024; Ding et al., 2024; Khare et al., 2023). While these studies only focused on method-level vulnerability detection, it reinforces our motivation that detecting vulnerabilities requires whole-project reasoning, which LLMs currently cannot do alone.
Static analysis tools. Apart from CodeQL (Avgustinov et al., 2016), other static analysis tools (CPPCheck, ; Semgrep, 2023; FlawFinder, ; FB Infer, ; Code Checker, ) also include analyses for vulnerability detection. More general query engines (Scholz et al., 2016; Li et al., 2023b) have also been applied to find program bugs. But these tools are not as feature-rich and effective as CodeQL (Li et al., 2023a; Lipp et al., 2022). Recently, proprietary tools such as Snyk (Snyk.io, ) and SonarQube (SonarQube, ) are also gaining in popularity, although sharing the same fundamental limitations of missing specifications and false positives, which IRIS improves upon. We envision our technique to benefit all such tools. Works such as MERLIN (Livshits et al., 2009), Seldon (Chibotaru et al., 2019) and InspectJS (Dutta et al., 2022) tackle the problem of specification inference through probabilistic modeling. Specifically, like IRIS, InspectJS also augments CodeQL with specifications inferred using machine learning. However, InspectJS relies on the quality of seed specifications and requires expensive analysis of each third-party library, which IRIS does not—making it more scalable. Future work could explore incorporating probability estimates for specifications.
LLM-based approaches for software engineering. Researchers are increasingly combining LLMs with program reasoning tools for challenging tasks such as fuzzing (Lemieux et al., 2023; Xia et al., 2024), program repair (Xia et al., 2023; Joshi et al., 2023; Xia & Zhang, 2022), and fault localization (Yang et al., 2023). While we are on a similar direction as (Li et al., 2024; Wang et al., 2024), to our knowledge, our work is among the first to combine LLMs with static analysis to detect application-level security vulnerabilities via whole-project analysis.
7 Conclusion and Limitations
We presented IRIS, a novel neuro-symbolic approach that combines LLMs with static analysis for vulnerability detection. We curate a dataset, CWE-Bench-Java, containing 120 security vulnerabilities across four classes in real-world projects. Our results show that systematically combining LLMs with static analysis significantly improves upon traditional static analysis alone in terms of both detected bugs and the alleviation of developer burden.
Limitations. There are still many vulnerabilities that IRIS cannot detect. Future approaches may explore a tighter integration of these two tools to improve performance. In addition, IRIS makes numerous calls to LLMs for specification inference and filtering false positives, increasing the potential cost of analysis. While our results on Java are promising, it is unknown if IRIS will perform well on other languages. Moreover, there is still a gap between the IRIS generated report and the report that the developers would like to see. We plan to explore this further in future work.
Acknowledgments
We thank the anonymous reviewers for their valuable feedback and suggestions that helped improve this work. We also thank Claire Wang for helping with open-sourcing the project. This research was supported by NSF award CCF 2313010.
Appendix A Implementation Details of IRIS
A.1 Selecting Candidate Specifications
While extracting external APIs, we filter out commonly-used Java libraries that are unlikely to contain any potential sources or sinks. Such libraries include testing libraries like JUnit and Hamcrest or mocking libraries like Mockito. While we filter out methods that are defined in the project, we specifically allow methods that are inherited from an external class or interface. An example is the getResource method of the generic class Class in java.lang package, which takes a path as a string and accesses a file in the module. Many projects commonly inherit this class and use this method. If the input path is unchecked, it may lead to a Path-Traversal vulnerability if the path accesses resources outside the given module. Hence, detecting such API usages is crucial.
Taint sources are typically values returned by methods that obtain inputs from external sources, such as response of an HTTP request or a command line argument. Hence, we select external APIs that have a “non-void” return type as candidate sources. Another type of taint sources are commonly seen in Java libraries. When used by downstream libraries, tainted information maybe passed into the library through function calls. Therefore, we also collect the formal parameters for public internal function as source candidates. Due to the excessive amount of such candidates, we pose a further constraint that the public internal function must be directly invoked by a unit test case within the same repository. Here, the test cases are identified by checking whether the residing file path has src/test within it.
On the other hand, taint sinks are typically arguments to an external API. This involves explicit arguments, such as the command argument passed to Runtime.exec(String command) method, and implicit this argument to non-static functions, such as the file variable in the function call file.delete(). This is the only type of sink that we consider within IRIS.
We note that this is not the entire story as there might be other kinds of sources and sinks. Other types of source candidates include the formal parameter of protected but overridden internal functions (the req parameter in protected HTTPServeletResponse doGet(HTTPServeletRequest req)), arguments to an impure external function (the buffer argument to void read(byte[] buffer, int size)), etc. Sink candidates include the return value of public facing functions, thrown exceptions, and even static methods without any parameter (System.exit()). Due to the complexity, we do not tackle such kind of sources of sinks in this work. However, we plan to explore further in future work.
A.2 LLM Prompts for Specification Inference
1System: You are a security expert. You are given a list of APIs to be labeled as potential taint sources, sinks, or APIs that propagate taints. Taint sources are values that an attacker can use for unauthorized and malicious operations when interacting with the system. Taint source APIs usually return strings or custom object types. Setter methods are typically NOT taint sources. Taint sinks are program points that can use tainted data in an unsafe way, which directly exposes vulnerability under attack. Taint propagators carry tainted information from input to the output without sanitization, and typically have non-primitive input and outputs. Return the result as a json list with each object in the format:
2
3{ "package": <package name>,
4 "class": <class name>,
5 "method": <method name>,
6 "signature": <signature of the method>,
7 "sink_args": <list of arguments or this‘; empty if the API is not sink>,
8 "type": <"source", "sink", or "taint-propagator"> }
9
10DO NOT OUTPUT ANYTHING OTHER THAN JSON.
11
12
13User: [CWE_LONG_DESCRIPTION]
14
15Some example source/sink/taint-propagator methods are:
16[CWE_SOURCE_SINK_EXAMPLES]
17
18Among the following methods, \
19assuming that the arguments passed to the given function is malicious, \
20what are the functions that are potential source, sink, or taint-propagators to [CWE_TITLE] attack (CWE-[CWE_ID])?
21
22Package,Class,Method,Signature
23[Package1],[Class1],[Method1],[Signature1]
24[Package2],[Class2],[Method2],[Signature2]
25[...]
Listing 1: LLM prompt for labelling external APIs as sources or sinks.
1System: You are a security expert. You are given a list of APIs implemented in established Java libraries, and you need to identify whether some of these APIs could be potentially invoked by downstream libraries with malicious end-user (not programmer) inputs. For instance, functions that deserialize or parse inputs might be used by downstream libraries and would need to add sanitization for malicious user inputs. On the other hand, functions like HTTP request handlers are typically final and wont be called by a downstream package. Utility functions that are not related to the primary purpose of the package should also be ignored. Return the result as a json list with each object in the format:
2
3{ "package": <package name>,
4 "class": <class name>,
5 "method": <method name>,
6 "signature": <signature>,
7 "tainted_input": <a list of argument names that are potentially tainted> }
8
9In the result list, only keep the functions that might be used by downstream libraries and is potentially invoked with malicious end-user inputs. Do not output anything other than JSON.
10
11
12User: You are analyzing the Java package [PROJECT_AUTHOR]/[PROJECT_NAME]. Here is the package summary:
13
14[PROJECT_README_SUMMARY]
15
16Please look at the following public methods in the library and their documentations (if present). What are the most important functions that look like can be invoked by a downstream Java package that is dependent on [PROJECT_NAME], and that the function can be called with potentially malicious end-user inputs? If the package does not seem to be a library, just return empty list as the result. Utility functions that are not related to the primary purpose of the package should also be ignored.
17
18Package,Class,Method,Doc
19[Package1],[Class1],[Method1],[Documentation1]
20[Package2],[Class2],[Method2],[Documentation2]
21[...]
Listing 2: LLM prompt for labeling formal parameters of internal APIs as sources.
[…]
The second prompt, depicted in Listing 2, is used to label the formal parameters of internal APIs as sources. Since we are analyzing internal API, the information such as project README and function documentations are commonly available. The goal is to find whether this internal API might be invoked by a downstream library with a malicious input passed to this formal parameter. This information is not CWE specific, hence no CWE information is included in this prompt.
We hypothesize that since LLMs are pre-trained on internet-scale data, they have knowledge about the behavior of widely used libraries and their APIs. Hence, it is natural to ask whether LLMs can be used to identify APIs that are relevant as sources or sinks for any vulnerability class. If successful, LLMs can alleviate manual effort, and drastically improve the effectiveness of static analysis tools.
A.3 Design Decisions of LLM Prompts
Here are some reasoning behind the design decisions for the LLM prompts:
  1. 1.
    Few-shot for labeling external API: Since the LLM is tasked with labeling source, sink, and taint-propagator APIs, we provide one example for each category, along with a negative example of an API that does not fall into any of these categories. We typically select examples from the Java standard libraries because they are widely used and their labels are readily available.
  2. 2.
    Zero-shot for labeling internal API: As labels for internal APIs are not available, we rely on the zero-shot capabilities of the language model. To mitigate potential performance loss, we include additional information, such as documentation associated with important internal APIs.
  3. 3.
    ±$\pm 5$ lines surrounding the source and sink location during contextual analysis: We chose ±$\pm 5$ lines as a balanced approach to provide sufficient context while managing performance and cost. While technically possible to use a larger window, we observed that excessive context can overwhelm the language model, leading to reduced accuracy. Additionally, a larger context increases computational costs significantly, particularly given the large number of candidate APIs and paths that must be queried.
  4. 4.
    Selecting subset of nodes in the alarm path during contextual analysis: We use a hyperparameter $S$ to control the number of intermediate steps included in the prompt. For paths with more than $S$ intermediate steps, we divide the path into $S$ equal segments and select one step from each. This selection prioritizes function calls, as they may indicate sanitizations. If no function call is present, a node is randomly selected from the segment. In our experiments, we observe that setting $S$ to 10 provides a good balance between the cost and accuracy so that the prompt contains enough context and would not be too long.
A.4 CodeQL Queries for Static Analysis
1import java
2// other imports ...
3import MySources
4import MySinks
5
6/**
7 * A taint-tracking configuration for tracking flow from remote sources to the
8 * creation of a path.
9 */
10module MyTaintedPathConfig implements DataFlow::ConfigSig {
11 predicate isSource(DataFlow::Node source) {
12 isLLMDetectedSource(source)
13 }
14
15 predicate isSink(DataFlow::Node sink) {
16 isLLMDetectedSink(sink)
17 }
18
19 predicate isBarrier(DataFlow::Node sanitizer) {
20 sanitizer.getType() instanceof BoxedType or
21 sanitizer.getType() instanceof PrimitiveType or
22 sanitizer.getType() instanceof NumberType or
23 sanitizer instanceof PathInjectionSanitizer
24 }
25
26 predicate isAdditionalFlowStep(DataFlow::Node n1, DataFlow::Node n2) {
27 isLLMDetectedStep(n1, n2)
28 }
29}
30
31/** Tracks flow from remote sources to the creation of a path. */
32module MyTaintedPathFlow = TaintTracking::Global<MyTaintedPathConfig>;
33
34from MyTaintedPathFlow::PathNode source, MyTaintedPathFlow::PathNode sink
35where MyTaintedPathFlow::flowPath(source, sink)
36select
37 getReportingNode(sink.getNode()),
38 source,
39 sink,
40 "This path depends on a \$@.",
41 source.getNode(),
42 sourceType(source.getNode())
Listing 3: CodeQL script for detecting vulnerabilities for Path-Traversal (CWE 22).
Listing 3 presents our CodeQL query for Path-Traversal vulnerabilities (CWE 22). Lines 10-29 describe a taint analysis configuration that describes which nodes in the data flow graph should be considered as sources or sinks. Here, Line 12 specifies our custom predicate isLLMDetectedSource that checks whether the method called is taint source based on our specifications. Similarly, our predicates isLLMDetectedSink checks whether the node is a taint sink based on our specifications. Line 16 checks if a method call or method argument node is a taint sink based on our specifications. We generate the source and sink specifications as predicates in QL file as shown in Listings 4 and 5 respectively. Given a taint configuration and the source and sink specifications, CodeQL can automatically perform taint analysis on a given project.
We use templates to convert LLM inferred specifications into CodeQL queries. There are three kinds of queries:
  1. 1.
    a formal parameter of an internal function as a source;
  2. 2.
    the return value of an external function as a source; and
  3. 3.
    an argument to an external function as a sink.
Example queries for the two kinds of sources are specified in Listing 4, while the example query for the sink is illustrated in Listing 5. As shown in the listings, we not only match on function package, class, and name, but also match on individual arguments or parameters. Moreover, our query handles generic functions or function in generic classes through the getSourceDeclaration() predicate provided by CodeQL. Notably, when the number of inferred specifications is too large, we will split the single predicate into multiple hierarchical ones, improving the CodeQL performance.
1predicate isLLMDetectedSource(DataFlow::Node src) {
2 // Sources: Return value from external APIs
3 (
4 src.asExpr().(Call).getCallee().getName() = "getName" and
5 src.asExpr().(Call).getCallee().getDeclaringType().getSourceDeclaration().hasQualifiedName("java.util.zip", "ZipEntry")
6 )
7 ...
8 or
9 // Sources: Function formal parameters of internal API
10 exists(Parameter p |
11 src.asParameter() = p and
12 p.getCallable().getName() = "setUserName" and
13 p.getCallable().getDeclaringType().getSourceDeclaration().hasQualifiedName("org.apache.dolphinscheduler.dao.entity", "DqRule") and
14 ( p.getName() = "userName" )
15 )
16 ...
17}
Listing 4: CodeQL predicate for source specifications.
1predicate isLLMDetectedSink(DataFlow::Node snk) {
2 exists(Call c |
3 c.getCallee().getName() = "createTempFile" and
4 c.getCallee().getDeclaringType().getSourceDeclaration().hasQualifiedName("java.io", "File") and
5 ( c.getArgument(0) = snk.asExpr().(Argument) )
6 )
7 or
8 ...
9}
Listing 5: CodeQL predicate for sink specifications.
A.5 Visualization of Metrics
We provide a visualization of our VulDetected metric in Fig. 10. For evaluation, we assume that the label for a project $P$ is provided as a set of crucial program points $\mathbf{V}_{\text{vul}}^{P}=\{V_{1},\dots,V_{n}\}$ where the vulnerable paths should pass through. In practice, these are typically the patched methods that can be collected from each vulnerability report. As illustrated in Fig. 10, if at least one detected vulnerable path passes through a fixed location for the given vulnerability, then we consider the vulnerability detected. Let $\textit{Paths}^{P}$ be the set of detected paths for each project $P$ from prior stages. The vulnerable paths inside project $P$ is given by:
$\begin{array}[]{rlrl}\textit{VulPaths}(P)=&\{\textit{Path}\in\textit{Paths}^{P}~{}|~{}\textit{Path}~{}\cap~{}\mathbf{V}_{\text{vul}}^{P}\neq\emptyset\}\end{array}$
Figure 10: A visualization of our metrics used for vulnerability detection, where the snippets are adapted from Zip4j v2.11.5-2, with slight changes for clearer presentation. While both sinks are potential causes of Path-Traversal (CWE-22), only the dataflow path on the left passes through the fixed sanitizer function. Therefore, we consider only the path on the left as a vulnerability.
Appendix B Additional Details of CWE-Bench-Java
B.1 Details of Dataset Extraction Process
Table 4: Vulnerability dataset collection statistics.
Step CWE-22 CWE-78 CWE-79 CWE-94 Total
Initial CVEs 236 39 681 109 1065
W/ Github URL and Version 119 37 219 55 430
W/ Fix Commit 89 27 99 50 265
Compilable 56 17 50 26 149
Fixes in Java Code 56 16 25 47 144
Manual Validation 55 13 31 21 120
Because we use CodeQL for static analysis, we further need to build each project for CodeQL to extract data flow graphs from the projects. To build each project, we need to determine the correct Java and Maven compiler versions. We developed a semi-automated script that tries to build each project with different combinations of Java and Maven versions. The fourth row in Table 4 presents the number of projects we were able to build successfully. Overall, this results in ($\star$) 149 projects.
Finally, we manually check each fix commit and validate whether the commit actually contains a fix to the given CVE in a Java file. For instance, we found that in some cases the fix is in files written in other languages (such as Scala or JSP). While code written in other languages may flow to the Java components in the project during runtime or via compilation, it is not possible to correctly determine if static analysis can correctly detect such a vulnerability. Hence, we exclude such CVEs. Further, we exclude cases where the vulnerability was in a dependency and the fix was just a version upgrade or if the vulnerability was mis-classified. Finally, we end up with ($\star$) 120 projects that we evaluate with IRIS. For this task, we divide the CVEs among two co-authors of the project, who independently validate each case. The co-authors cross-check each other’s results and discuss together to come up with the final list of projects.
The closest dataset to ours, in terms of features, is the Java dataset curated by Li et al. (2023a), containing 165 CVEs. While we initially considered using their dataset for our work, we found several issues. First, their dataset does not come with build scripts, which makes it hard to automatically run each project with CodeQL. Second, their dataset only has few CVEs for all but one CWE, which makes it difficult to thoroughly analyze a tool for different vulnerability classes. Finally, they do not provide any automated scripts to curate more CVEs. Hence, we curated our own dataset and our framework also allows to easily extend to more vulnerability classes.
B.2 Comparison of our CWE-Bench-Java with Existing Vulnerability Datasets
We compare CWE-Bench-Java with existing datasets for vulnerability detection in Java, C, and C++ codebases, on the following criteria:
  1. 1.
    CVE Info: whether CVE Metadata is contained in the dataset;
  2. 2.
    Real-World: whether the dataset contains real-world projects;
  3. 3.
    Fix Locations: whether the dataset contains fix information about the vulnerabilities;
  4. 4.
    Compilable: whether the dataset ensures that the projects are end-to-end and automatically compilable; and
  5. 5.
    Vetted: whether the vulnerability in the dataset is manually verified and confirmed.
As shown in Table 5, compared to existing datsets, CWE-Bench-Java, is the only one that checks every criterion. This underscores the significance of our new dataset.
Table 5: Comparison of CWE-Bench-Java with existing vulnerability datasets.
Dataset Languages CVE Info Real-World Fix Locations Compilable Vetted
BigVul C/C++
Reveal C/C++
CVEFixes C/C++, Java, …
DiverseVul C/C++
DeepVD C/C++
Juliet C++, Java
Li et al. (2023a) Java
SVEN (He & Vechev, 2023) C++
CWE-Bench-Java (Ours) Java
Appendix C Evaluation Details
C.1 Experimental Settings
We select two closed-source LLMs from OpenAI: GPT 4 (gpt-4-0125-preview) and GPT 3.5 (gpt-3.5-turbo-0125) for our evaluation. GPT 4 and GPT 3.5 queries used in the paper are performed through OpenAI API during April and May of 2024.
We also select instruction-tuned versions of six state-of-the-art open-source LLMs via huggingface API: Llama 3 8B and 70B, DeepSeekCoder 7B, QWEN-2.5-Coder 32B, and Gemma-2 27B. To run the open-source LLMs we use two groups of machines: a 2.50GHz Intel Xeon machine, with 40 CPUs, four GeForce RTX 2080 Ti GPUs, and 750GB RAM, and another 3.00GHz Intel Xeon machine with 48 CPUs, 8 A100s, and 1.5T RAM.
We use CodeQL version 2.15.3 as the backbone of our static analysis. We have patched CodeQL with an additional feature that augments the Dataflow edge between throw statement and its closest surrounding try-catch block. We use this CodeQL pull request as the base of our patch.
C.2 CodeQL Baseline
For baseline comparison with CodeQL, we use the built-in Security queries specifically designed for each CWE that comes with CodeQL 2.15.3. Note that there are multiple security queries for each CWE, and each produce alarms of different levels (error, warning, and recommendation). For each CWE, we take the union of alerts generated by all queries and do not differentiate between alarms of different levels. For instance, there are 3 queries from CodeQL for detecting CWE-22 vulnerabilities, namely TaintedPath, TaintedPathLocal, and ZipSlip. While TaintedPath and ZipSlip produce error level alarms, TaintedPathLocal produces only alarm recommendations. To CodeQL’s advantage, all alarms are treated equally in our comparisons.
C.3 Hyper-Parameters and Few-Shot Examples
During IRIS, we have 2 prompts that are used to label external and internal APIs. Recall that the prompts contain batched APIs. We use batch size of 20 and 30 for internal and external, respectively. In terms of few-shot examples passed to labeling external APIs, we use 4 examples for CWE-22, 3 examples for CWE-78, 3 examples for CWE-79, and 3 examples for CWE-94. We use a temperature of 0, maximum tokens to 2048, and top-p of 1 for inference with all the LLMs. For GPT 3.5 and GPT 4, we also fix a seed to mitigate randomness as much as possible.
C.4 Details of Selected LLMs
We include the versions of the 7 selected LLMs in Table 6.
Table 6: Versions and model IDs of the selected LLMs in our evaluation.
LLM Version and Size Model ID
GPT 4 gpt-4-0125-preview
GPT 3.5 gpt-3.5-turbo-0125
Llama 3 8B meta-llama/Meta-Llama-3-8B-Instruct
Llama 3 70B meta-llama/Meta-Llama-3-70B-Instruct
DeepSeekCoder 7B deepseek-ai/deepseek-coder-7b-instruct
Qwen-2.5-Coder 32B Instruct Qwen/Qwen2.5-Coder-32B-Instruct
Gemma-2 27B google/gemma-2-27b-it
C.5 Statistics of Unique and Recurring Specifications
Table 7: Unique source and sink specifications across all projects in CWE-Bench-Java.
CWE 22 78 79 94
#Unique Sources 1348 899 598 810
#Unique Sinks 1069 575 514 1281
Table 8: Recurring source and sink specifications in CWE-Bench-Java.
CWE 22 78 79 94
#Recurring Sources 908 232 1118 626
#Recurring Sinks 919 201 911 961
Continuous taint specification inference is necessary. Our results show that there is a high number of both unique and recurring sources and sinks. Table 8 presents the number of inferred source and sink specifications that occur only in a single project in CWE-Bench-Java, whereas Table 8 presents the specifications that occur in at least two projects. This indicates that even if previously inferred specifications are useful, a significant number of new relevant APIs still remain and need to be labeled for effective vulnerability detection. This observation strongly motivates the design of IRIS that infers these specifications on-the-fly for each project via LLMs, instead of relying on a fixed corpus of specifications like CodeQL.
C.6 Statistics of Inferred Taint Specifications
We show the statistics of inferred taint specifications in Table 9. As shown by the percentage, GPT-4 generates smaller set of sources and sinks than smaller-scale LLMs like DeepSeekCoder 7B.
Table 9: Ratio of API candidates labeled as source (S) or sink (N) by GPT-4 and DeepSeekCoder (DSC) 7B, per CWE and in total.
CWE #Cand. GPT-4 DSC 7B
%S %N %S %N
22 130,974 2.03% 1.90% 4.27% 4.01%
78 25,605 4.73% 3.37% 3.67% 3.33%
79 37,138 5.69% 4.69% 4.28% 4.56%
94 36,325 5.12% 7.83% 6.11% 6.21%
Total 230,042 3.41% 3.45% 4.50% 4.37%
C.7 Error Analysis: Cause of Undetected Vulnerabilities
In general, we find the following three main causes of undetected vulnerabilities by employing IRIS:
  1. 1.
    Vulnerability cannot be modeled by simple taint dataflows: for instance, the vulnerability CVE-2020-11977 (CWE-94) is manifested by an unexpected exit(1) call, with no direct taint dataflow going into it. In fact, the taint dataflow goes into the condition of an “if” statement surrounding the exit. In this case, the model of sink API cannot capture the vulnerability.
  2. 2.
    Missing dataflow edge due to side-effects: for instance a taint information is written to a temporary file, and is later read from the same file, subsequently causing a vulnerability. However, the dataflow edge is carried through side-effects, which is not captured by CodeQL.
  3. 3.
    Missing dataflow edge due to unspecified library usage: The vulnerability can only be manifested through a concrete usage of the library; but within the library itself there is no possible dataflow to connect the source and the sink.
Overall, we view the above as general limitations for the static analysis. In terms of LLM induced false negatives, here are the two main failure modes:
  1. 1.
    Missing taint-propagator labels from LLM: this would cause missing dataflow edges stopping source to flow to sink.
  2. 2.
    Missing source or sink specifications from LLM: if there is no relevant source/sink specification then the static analysis tool would not have the anchor for analysis.
Appendix D Analysis Runtime
We include the full table containing statistics to provide more details about projects and our analysis (Table LABEL:tab:longtable). For each project, we present its corresponding CWE ID, the lines-of-code (SLOC), the time it takes to run the full analysis, the number candidate APIs and the number of labeled source and sinks by Llama 3 8B. We also color code cells of interest: For SLOC, we mark a cell as red if $>$1M; yellow if $>$100k. For Time, we mark a cell as red if $\geq$1h; yellow if $\geq$5m. For the number of candidates, we mark a cell as red if $>$10k. Lastly for the numbers of sources and sinks, we mark a cell as red if the number is larger than 200.
Table 10: Details of analysis runtime, candidates, and inferred sources and sinks for all projects (Llama 3 8B).
CWE-ID Project SLOC Time #Candidates #Sources #Sinks
22 DSpace 218.2K 15s 3.61K 162 217
22 spark 10.7K 1m 679 35 27
22 spark 9.77K 57s 598 33 22
22 wildfly 496.28K 4m 14.13K 457 425
22 vertx-web 51.01K 1m 2.06K 80 77
22 camel 1.16M 8m 293 22 9
22 hutool 135.34K 4m 6.17K 115 211
22 tika 106.3K 2m 3.84K 277 177
22 retrofit 19.28K 1m 880 28 13
22 jspwiki 149.45K 1m 1.83K 62 80
22 camel 1.21M 11m 4.43K 53 80
22 tapestry-5 160.06K 1m 3.04K 91 66
22 spring-cloud-co 18.56K 1m 1.16K 40 64
22 spring-cloud-co 18.44K 59s 1.16K 40 64
22 rocketmq 94.64K 1m 2.78K 28 54
22 mpxj 181.55K 1m 1.6K 37 43
22 flink 1.14M 2h 5.16K 39 61
22 java 1M 2m 8.04K 96 41
22 commons-io 29.24K 58s 1.07K 12 47
22 karaf 135.22K 1m 5.43K 150 210
22 james-project 434.32K 4m 14.58K 209 226
22 vertx-web 49.28K 1m 2.36K 83 96
22 esapi-java-lega 35.26K 59s 1.48K 43 67
22 xwiki-commons 103.05K 1m 3.76K 104 137
22 zip4j 16.78K 58s 532 6 34
22 one-java-agent 5.19K 51s 327 11 20
22 myfaces 161.02K 1m 2.4K 68 44
22 undertow 86.03K 1m 2.58K 66 93
22 DependencyCheck 28.57K 1m 1.23K 47 66
22 plexus-archiver 13.04K 51s 573 34 47
22 plexus-archiver 13.04K 51s 573 34 47
22 zt-zip 6.64K 52s 337 14 31
22 curekit 511 43s 73 2 4
22 aws-sdk-java 7.72M 38m 12K 62 65
22 venice 115.44K 1m 2.27K 36 79
22 DSpace 237.33K 1m 3.67K 179 233
22 Payara 1.12M 7m 16.05K 379 427
22 DSpace 237.33K 1m 3.67K 179 233
22 goomph 12.68K 59s 1.12K 35 111
22 dolphinschedule 90.69K 1m 3.36K 65 92
22 dolphinschedule 91.94K 1m 3.4K 65 92
22 testng 95.53K 1m 2.08K 33 73
22 uima-uimaj 226.81K 2m 5.66K 103 176
22 keycloak 614.82K 12m 13.34K 325 252
22 glassfish 1.19M 5m 12.19K 293 346
22 graylog2-server 382K 4m 13.3K 227 171
22 mina-sshd 130.14K 1m 3.64K 52 120
22 shiro 38.68K 1m 1.5K 41 42
22 plexus-archiver 15.51K 57s 666 37 56
22 plexus-utils 23.3K 58s 754 16 36
22 yamcs 693.6K 2m 11K 98 113
22 yamcs 693.6K 2m 11K 98 113
22 shiro 38.94K 1m 1.53K 41 43
22 sling-org-apach 8.34K 54s 695 28 25
78 xstream 43.49K 1m 1.39K 91 30
78 xstream 59.79K 1m 1.64K 107 42
78 xstream 52.25K 1m 1.64K 107 43
78 docker-commons- 2.79K 54s 362 25 20
78 workflow-cps-pl 17.02K 1m 1.38K 72 61
78 workflow-cps-gl 4.31K 55s 523 40 38
78 workflow-multib 3.45K 53s 500 30 30
78 activemq 442.42K 4m 6.34K 234 192
78 plexus-utils 22.76K 1m 714 34 17
78 git-client-plug 16.41K 1m 1.06K 83 50
78 perfecto-plugin 667 54s 107 5 10
78 nifi 915.95K 11m 22.44K 894 614
78 script-security 8.17K 1m 678 40 46
79 antisamy 6.38K 57s 381 42 33
79 antisamy 6.38K 56s 381 42 33
79 jspwiki 149.33K 1m 1.84K 156 110
79 jspwiki 149.33K 1m 1.84K 156 110
79 jspwiki 149.33K 1m 1.84K 156 110
79 jspwiki 157.09K 1m 1.85K 157 110
79 hibernate-valid 93.6K 1m 2.06K 79 57
79 cxf 798.53K 1h 16.54K 821 756
79 xxl-job 9.32K 60s 540 42 41
79 json-sanitizer 1.47K 52s 67 4 5
79 hawkbit 112.09K 1m 4.07K 144 151
79 nacos 203.78K 2m 4.08K 201 139
79 antisamy 4.93K 1m 362 43 34
79 esapi-java-lega 35.26K 1m 1.48K 107 85
79 antisamy 5.14K 1m 377 44 36
79 jolokia 29.97K 1m 1.66K 117 97
79 keycloak 60.6K 1m 2.1K 170 136
79 cxf 722.83K 15m 15.09K 766 710
79 sling-org-apach 1.37K 55s 136 4 13
79 DSpace 237.33K 2m 3.67K 347 320
79 keycloak 615.6K 3h 13.37K 606 461
79 keycloak 615.6K 3h 13.37K 606 461
79 xwiki-commons 105.92K 1m 3.94K 244 151
79 xwiki-commons 105.94K 1m 3.94K 244 151
79 xwiki-rendering 97.01K 1m 1.22K 73 92
79 xwiki-commons 106.87K 1m 3.99K 254 161
79 jspwiki 158.7K 1m 2.22K 176 126
79 keycloak 617.15K 4h 14.04K 643 479
79 xwiki-commons 107.09K 1m 3.03K 209 143
79 jstachio 53.02K 54s 792 40 46
79 xwiki-rendering 97.63K 1m 1.24K 74 92
94 spring-security 43.9K 1m 1.83K 120 176
94 xstream 52.25K 1m 1.64K 111 145
94 cron-utils 13.08K 1m 476 13 26
94 struts 160.51K 12m 4.39K 301 357
94 activemq 547.68K 1h 7.55K 370 607
94 spring-framewor 666.11K 45m 17.71K 688 846
94 spring-cloud-ga 25.56K 1m 2.01K 130 153
94 dubbo 175.63K 2m 6.73K 342 383
94 incubator-dubbo 96.35K 1m 3.68K 194 255
94 spring-security 57.34K 1m 2.43K 192 234
94 kubernetes-clie 806.35K 3m 2.33K 93 130
94 commons-text 24.87K 1m 962 40 47
94 ff4j 46.21K 1m 2.39K 133 274
94 spring-boot-adm 18.29K 1m 1.83K 92 157
94 sqlite-jdbc 17.71K 59s 732 50 74
94 nifi 993.76K 25m 57 2 11
94 rocketmq 108.39K 2m 3.4K 117 164
94 nifi 1.01M 27m 261 27 24
94 rocketmq 197.78K 2m 6.28K 205 252
94 dolphinschedule 154.95K 4m 5.78K 229 353
94 dolphinschedule 154.95K 4m 5.78K 229 353